OpenStudy (anonymous):

how many solutions are there in the equation below. 5x-8(x+5)=6-3x

OpenStudy (gorv):

open the bracket

OpenStudy (gorv):

5x-8x-40=6-3x -3x-40=6-3x

OpenStudy (gorv):

-40=6 not possible so no solution